OpenStudy (neo92):

dy/dx = 5y , with y(0) = 3 acan any one here tel me how solve this

jimthompson5910 (jim_thompson5910):

dy/dx = 5y dy/dx = 5y dy = 5y*dx dy/y = 5dx Integrate both sides and tell me what you get

OpenStudy (neo92):

\[\ln y = 5x\]

jimthompson5910 (jim_thompson5910):

more like ln(|y|) = 5x + C but you pretty much have it Now solve for y

OpenStudy (neo92):

yeah.thnx actually thats ware im stucked.. is it \[y=e ^{(5x+C)}\]

jimthompson5910 (jim_thompson5910):

ln(|y|) = 5x + C |y| = e^(5x + C) |y| = e^(5x)*e^(C) |y| = e^(C)*e^(5x) |y| = A*e^(5x) y = +-A*e^(5x) where A = e^C

jimthompson5910 (jim_thompson5910):

now use the idea that y(0) = 3

jimthompson5910 (jim_thompson5910):

to find the value of 'A' (which will help you find C, but that's a bit irrelevant)

OpenStudy (neo92):

thnx... got it solved :D

jimthompson5910 (jim_thompson5910):

yw